OpenStudy (anonymous):

Ask a question... Find the coordinates of the intersection of the diagnols of the parallelograms XYZW with vertices X (2,2) Y(3,6) Z(10,6) and W(9,2)

OpenStudy (anonymous):

THERE IS NO ANSWER\[\omega \pi \theta \rho \iota \kappa \iota \omega \psi \theta \iota \kappa \theta \xi \chi \theta \pi\]